Melanie can paint a certain barn by herself in x days . Her helper Maleissa can paint the same barn by her self in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together . Evaluate the expression for x=5.
Answer by stanbon(75887)

Melanie can paint a certain barn by herself in x days . Her helper Maleissa can paint the same barn by her self in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together . Evaluate the expression for x=5.
Melanie DATA:
time= x days/job ; rate= 1/x job/day
Maleissa DATA:
time= 2x days/job ; rate = 1/2x job/day
Together DATA:
rate working together = 1/x + 1/2x = 3/2x job/day
Evaluate the expression for x=5:
3/10 job/day
Cheers,
Stan H.